# How Can Time Series Analysis Benefit From Multiple Modalities? A Survey and Outlook

### v1

@@ 2. Background and Taxonomy
(Alcaraz and Strodthoff, 2023; Chung et al., 2023; Lai et al., 2025; Zhao et al., 2024b; Guo et al., 2024; Yu et al., 2024; Liu et al., 2023c; Chan et al., 2024; Cosentino et al., 2024), fill=purple!0] ] [Financial TS, fill=red!10 [(Liu et al., 2022a, 2023a; Alcaraz and Strodthoff, 2023; Chen et al., 2024b; Li et al., 2023a; Bhatia et al., 2024; Xie et al., 2024c), fill=purple!0] ] ] [Text 2 Time + Time 2 Text (§5.3), fill=green!20 [(Wang et al., 2024b; Xie et al., 2024b; Kong et al., 2025; Kim et al., 2024b; Islam et al., 2024) ,fill=purple!0] ] [Text 2 Time (§5.1), fill=green!20 [Generation, fill=orange!10 [(Lai et al., 2025; Merrill et al., 2024; Xie et al., 2024b), fill=purple!0] ] [Retrieval, fill=orange!10 [(Fons et al., 2024; Cai et al., 2024; Ito et al., 2024), fill=purple!0] ] ] [Time 2 Text (§5.2), fill=green!20 [Explanation, fill=orange!10 [(Aksu et al., 2024; Jiang et al., 2025; Liu et al., 2024i), fill=purple!0] ] [Captioning, fill=orange!10 [(Sharma et al., 2021; Spreafico and Carenini, 2020; Li et al., 2023b; Zhang et al., 2023b; Dohi et al., 2024; Lin et al., 2024; Trabelsi et al., 2025), fill=purple!0] ] ] ] [Time+X:
@@
(Liu et al., 2024d; Darapaneni et al., 2022; Tavakoli et al., 2025; Rekabsaz et al., 2017; Bamford et al., 2023; Dong et al., 2024), fill=purple!0] ] ] [Time + Other Modalities (§4.2), fill=green!20 [(Ebrahimi et al., 2023; Girdhar et al., 2023), fill=purple!0] ] [Time + Text (§4.1), fill=green!20 [(Ming et al., 2023; Liu et al., 2024j; Wei et al., 2022; Williams et al., 2024; Jin et al., 2023; Wu et al., [n. d.]; Wang et al., 2024b; Liu et al., 2025a; Xu et al., 2024a; Dong et al., 2024; Liu et al., 2024b; Xinlei et al., 2024; Cheng et al., 2024; Tao et al., 2024; Liu et al., 2024f; Li et al., 2025b; Zhang et al., 2024c), fill=purple!0] ] ] [TimeAsX:
@@
(Li et al., 2022; Bao et al., 2022; Jiang et al., 2023; Tang et al., 2024a; Ren et al., 2024), fill=purple!0] ] [Medical TS, fill=red!10 [(Belyaeva et al., 2023; Liu et al., 2023d, 2024a; Zhao et al., 2024b; Zhang et al., 2024c; Chan et al., 2024), fill=purple!0] ] [Financial TS, fill=red!10 [(Xu et al., 2024b; Sezer and Ozbayoglu, 2019; Pei et al., 2024; Zeng et al., 2023; Dosovitskiy et al., 2020; Du et al., 2020)
(Chen and Tsai, 2020; Li et al., 2020; Wang and Oates, 2015; Hu et al., 2024; Su et al., 2024; Foroutan and Lahmiri, 2024; Cheng et al., 2022), fill=purple!0] ] ] [Time As Other Modalities (§3.2), fill=green!20 [Table, fill=green!20, [(Inkit et al., 2020; Dooley et al., 2023; Hoo et al., [n. d.]; Wang et al., 2024a), fill=purple!0] ] [Audio, fill=green!20, [(Chao-Han et al., 2021; Gong et al., 2022; Minhao et al., 2021) ,fill=purple!0] ] ] [Time As Image (§3.2), fill=green!20 [(Zekun et al., 2023; Luoxiao et al., 2024; Zhuang et al., 2024; Xu et al., 2025; Yang et al., 2024; Daswani et al., 2024; Mouxiang et al., 2024; Dosovitskiy et al., 2020; Zeng et al., 2021; Namura and Ichikawa, 2024; Roth et al., 2022; Zhiguang and Tim, 2015; Kim et al., 2024a; Li et al., 2020; Siru et al., 2025), fill=purple!0] ] [Time As Text (§3.1), fill=green!20 [(Gruver et al., 2023; Liu et al., 2024j; Xinglei et al., 2023; Hao et al., 2022; Zhou et al., 2023; Rasul et al., 2023; Yuxuan et al., 2024; Ching et al., 2023; Chenxi et al., 2023; Cao et al., 2024; Ming et al., 2023; Liu et al., 2024e)
@@ 5.5. Domain-Specific Time-Series Works
FinTime 2 Text. This type of cross-modality work for finance domain applications mainly focuses on (1) Financial Report Generation, converting time series data combined with other data types such as tables into readable financial reports, incorporating statistical models, market analysis, and risk assessment metrics (Liu et al., 2022a, 2023a; Alcaraz and Strodthoff, 2023) (2) Market Analysis and Sentiment Reports, correlating economic indicators, market sentiment, and time series data to generate market interpretation text (Chen et al., 2024b; Li et al., 2023a). FinTime 2 Text + Text 2 FinTime. This type of work builds foundation finance chat models (Bhatia et al., 2024; Xie et al., 2024c).

### v2

@@ 1. Introduction
(Alcaraz and Strodthoff, 2023; Chung et al., 2023; Lai et al., 2025; Zhao et al., 2024b; Guo et al., 2024; Yu et al., 2024; Liu et al., 2023c; Chan et al., 2024; Cosentino et al., 2024), fill=purple!0] ] [Financial TS, fill=red!10 [(Liu et al., 2022a, 2023a; Chen et al., 2024b; Li et al., 2023a; Bhatia et al., 2024; Xie et al., 2024c), fill=purple!0] ] ] [ Other Cross-Modality Works (§5.4, fill=green!20 [(Islam et al., 2024) ,fill=purple!0] ] [Text 2 Time + Time 2 Text (§5.3), fill=green!20 [(Wang et al., 2024b; Xie et al., 2024b; Kong et al., 2025; Kim et al., 2024b) ,fill=purple!0] ] [Text 2 Time (§5.1), fill=green!20 [Generation, fill=orange!10 [(Lai et al., 2025; Merrill et al., 2024; Xie et al., 2024b), fill=purple!0] ] [Retrieval, fill=orange!10 [(Fons et al., 2024; Cai et al., 2024; Ito et al., 2024), fill=purple!0] ] ] [Time 2 Text (§5.2), fill=green!20 [Explanation, fill=orange!10 [(Aksu et al., 2024; Jiang et al., 2025; Liu et al., 2024i), fill=purple!0] ] [Captioning, fill=orange!10 [(Sharma et al., 2021; Spreafico and Carenini, 2020; Li et al., 2023b; Zhang et al., 2023b; Dohi et al., 2024; Lin et al., 2024; Trabelsi et al., 2025), fill=purple!0] ] ] ] [Time+X:
@@
(Liu et al., 2024d; Darapaneni et al., 2022; Tavakoli et al., 2025; Rekabsaz et al., 2017; Bamford et al., 2023; Dong et al., 2024), fill=purple!0] ] ] [Time + Other Modalities (§4.2), fill=green!20 [(Ebrahimi et al., 2023; Girdhar et al., 2023), fill=purple!0] ] [Time + Text (§4.1), fill=green!20 [(Ming et al., 2023; Liu et al., 2024j; Williams et al., 2024; Wu et al., [n. d.]; Wang et al., 2024b; Liu et al., 2025a; Xu et al., 2024a; Xinlei et al., 2024; Cheng et al., 2024; Tao et al., 2024; Liu et al., 2024f; Li et al., 2025b; Zhang et al., 2024c), fill=purple!0] ] ] [TimeAsX:
@@
(Li et al., 2022; Bao et al., 2022; Jiang et al., 2023; Tang et al., 2024a; Ren et al., 2024; Xinglei et al., 2023; Hao et al., 2022), fill=purple!0] ] [Medical TS, fill=red!10 [(Belyaeva et al., 2023; Liu et al., 2023d, 2024a; Zhao et al., 2024b; Zhang et al., 2024c; Chan et al., 2024), fill=purple!0] ] [Financial TS, fill=red!10 [(Xu et al., 2024b; Sezer and Ozbayoglu, 2019; Pei et al., 2024; Zeng et al., 2023; Dosovitskiy et al., 2020; Du et al., 2020)
(Chen and Tsai, 2020; Li et al., 2020; Wang and Oates, 2015; Hu et al., 2024; Su et al., 2024; Foroutan and Lahmiri, 2024; Cheng et al., 2022), fill=purple!0] ] ] [Time As Other Modalities (§3.2), fill=green!20 [Table, fill=green!20, [(Inkit et al., 2020; Dooley et al., 2023; Hoo et al., [n. d.]; Wang et al., 2024a), fill=purple!0] ] [Audio, fill=green!20, [(Chao-Han et al., 2021; Gong et al., 2022; Minhao et al., 2021) ,fill=purple!0] ] ] [Time As Image (§3.2), fill=green!20 [(Zekun et al., 2023; Luoxiao et al., 2024; Zhuang et al., 2024; Xu et al., 2025; Yang et al., 2024; Daswani et al., 2024; Mouxiang et al., 2024; Dosovitskiy et al., 2020; Zeng et al., 2021; Namura and Ichikawa, 2024; Roth et al., 2022; Zhiguang and Tim, 2015; Kim et al., 2024a; Li et al., 2020; Siru et al., 2025; Liu et al., 2024c), fill=purple!0] ] [Time As Text (§3.1), fill=green!20 [(Gruver et al., 2023; Liu et al., 2024j; Zhou et al., 2023; Rasul et al., 2023; Yuxuan et al., 2024; Ching et al., 2023; Chenxi et al., 2023; Cao et al., 2024; Ming et al., 2023; Liu et al., 2024e)
@@ 5.5. Domain-Specific Time-Series Works
FinTime 2 Text. This type of cross-modality work for finance domain applications mainly focuses on (1) Financial Report Generation, converting time series data combined with other data types such as tables into readable financial reports, incorporating statistical models, market analysis, and risk assessment metrics (Liu et al., 2022a, 2023a) (2) Market Analysis and Sentiment Reports, correlating economic indicators, market sentiment, and time series data to generate market interpretation text (Chen et al., 2024b; Li et al., 2023a). FinTime 2 Text + Text 2 FinTime. This type of work builds foundation finance chat models (Bhatia et al., 2024; Xie et al., 2024c).